# Quantum automata and languages of finite index

### v2

@@ 6.2 The case of monoidal languages
By Lemma 9, for every $A\in V_{1}$, there exists an effective computable regular subset $\cal R$ of $\ \Sigma_{1}^{*}\times\Sigma_{1}^{*}$ such that $M_{A}=\zeta({\cal R})$. We will prove the following slightly more general claim:
Claim. For every regular subset $\cal R$ of $\Sigma_{1}^{*}\times\Sigma_{1}^{*}$, ${\mathbf{C}l}(\zeta({\cal R}))$ contains $I$ and ${\mathbf{C}l}(\zeta({\cal R}))=\Psi(H)$ for some $H\in({\cal H})$ (cf (33)) and regular map $\Psi$. Further, if ${\cal R}=R^{*}$, then ${\mathbf{C}l}(\zeta({\cal R}))\in({\cal H})$.
We first show the claim for a finite set $\cal R$. We may suppose $\cal R$ =$=\{r\}$, where $r=(\alpha,\beta)$ with $\alpha=A_{1}A_{2}\cdots A_{s},$ $\beta=B_{1}\cdots B_{t}$, $A_{i},B_{i}\in\Sigma_{1},\ s,t\geq 0$. We assume fist that $\alpha$ and $\beta$ have the same length, i.e., $s=t$. By (36), the fact that $\zeta$ is a monoid morphism and Corollary 1, one gets
@@
This proves that the claim for $\cal R$ =$=\{r\}$. Finally, if $s>t$, then we can write the word $\beta=B_{1}\cdots B_{t}$ as $\beta=B_{1}\cdots B_{s}$, with $B_{j}=\varepsilon$, for $j=t+1,\ldots,s$. Afterwards, set $L_{B_{j}}=\{\varepsilon\}$ and $\varphi(L_{B_{j}})=\{I\}$, $t+1\leq j\leq s$. Then one proceeds to prove the claim, by repeating the argument above for $s=t$.
We now prove the claim according to the following cases:
@@
where
${\mathbf{C}l}(\zeta({\cal R}_{1}))=\Psi_{1}(H_{1}),\quad{\mathbf{C}l}(\zeta({\cal R}_{2}))=\Psi_{2}(H_{2}),$ (41)
with $H_{1},H_{2}\in({\cal H})$ and $\Psi_{1},\Psi_{2}$ being regular maps.
(i) By (41) one gets ${\mathbf{C}l}(\zeta({\cal R}))=\Psi(H_{1}\cup H_{2})$, and the claim comes from $H_{1}\cup H_{2}\in({\cal H})$.
(ii) By Corollary 1 and the fact that $\zeta$ is a monoid morphism, (41) yields ${\mathbf{C}l}(\zeta({\cal R}))=\Psi_{1}(H_{1})\Psi_{2}(H_{2}).$ From this, since $I\in\Psi_{1}(H_{1}),\Psi_{2}(H_{2})$ then $I\in{\mathbf{C}l}(\zeta({\cal R}))$. Now, observe that, by Propositions 1, 4, and 5, the condition $H_{1},H_{2}\in({\cal H})$ implies $H_{1}\oplus H_{2}\in({\cal H})$. The claim follows from the fact that
$\Psi_{1}(H_{1})\Psi_{2}(H_{2})=\Psi(H_{1}\oplus H_{2}),$
where $\Psi$ is the regular map defined as $\Psi(X\oplus Y):=\Psi_{1}(X)\Psi_{2}(Y)$.
(iii) Since ${\cal R}=R^{*}$ and the fact that $\zeta$ is a monoid morphism, then one has
${\mathbf{C}l}(\zeta({\cal R}))={\mathbf{C}l}(\zeta({R^{*}}))={\mathbf{C}l}(\zeta({R})^{*}).$
Observe that, from this, it trivially follows that ${\mathbf{C}l}(\zeta({\cal R}))$ contains $I$.
@@
one has
${\mathbf{C}l}(\zeta({\cal R}))={\mathbf{C}l}({\mathbf{C}l}(\zeta({R}))^{*}).$
By Theorems 2 and 3, one has
${\mathbf{C}l}({\mathbf{C}l}(\zeta({R}))^{*})=\overline{{\mathbf{C}l}(\zeta({R}))^{*}},$ (43)
where $\overline{{\mathbf{C}l}(\zeta({R}))^{*}}$ is the Zariski closure of ${\mathbf{C}l}(\zeta({R})^{*})$. Recall that, by (41), one has ${\mathbf{C}l}(\zeta({R}))=\Psi(H)$, where $\Psi$ is a regular map and $H\in({\cal H})$. This fact and (43) now implies
${\mathbf{C}l}(\zeta({\cal R}))=\overline{\Psi(H)^{*}}.$
Since $H\in({\cal H})$, we can write $H=\bigcup_{i=1}^{k}\ H_{i}$ as a finite union of irreducible effective algebraic sets $H_{i}$, every one of which containing the identity matrix $I$. This, together with Lemma 2, implies
${\mathbf{C}l}(\zeta({\cal R}))\ =\ \overline{(\Psi(H_{1})\ \cdots\ \Psi(H_{k}))^{*}}.$
Observe that in the formula above, we can write $\Psi(H_{1})\ \cdots\ \Psi(H_{k})$ as
$\widehat{\Psi}(H_{1}\oplus\cdots\oplus H_{k})$, where $\widehat{\Psi}$ is the regular map defined as $\widehat{\Psi}(X_{1}\oplus\cdots\oplus X_{k})=\Psi(X_{1})\cdots\Psi(X_{k})$. Hence we get
${\mathbf{C}l}(\zeta({\cal R}))\ =\overline{(\widehat{\Psi}(H))^{*}},$
where, again, $H=(H_{1}\oplus\cdots\oplus H_{k})\in({\cal H})$ as one can easily checks. The claim follows by applying Proposition 6. This completes the proof. ∎∎
@@
Proof of Proposition 11: As remarked before, it is enough to prove the claim for $k\leq 2$ since, by (17), the general case is treated with the same argument. Let $G=\langle V,\Sigma,P,S\rangle$ and observe that, for every $A\in V$, ${\mathbf{C}l}(M_{A})$ is effective algebraic by hypothesis if $k=1$ and, by Lemma 10, if $k=2$. The claim now follows from Proposition 9. ∎
@@
An example describing the main steps of Proposition 11 will be given in Section 8.
@@ 7 Concluding remarks and open problems
[$\alpha$]  One could extend Proposition 11 to a broader class of quantum automata. In this context, the construction of the decidability procedure would require the computation of the Zariski closure of a group $G$ of matrices that is algebraically presented, i.e. $G$ admits a set of generators that is algebraic itself. It may be of interest to investigate conditions ensuring the extension of the algorithms of [12, 17, 25] for such groups.
@@
[$\gamma$]  The $(L,Q)$ Intersection problem is decidable for bounded semi-linear languages. It may be of interest to get extensions of this result for the broader class of languages accepted by the reversal bounded non deterministic counter machines [18]. This extension could exploit a characterization of such languages provided by Ibarra in [19] in terms of some special combinatorial systems called RLGMC-grammars (right-linear grammar with multi-counters). Essentially, a RLGMC-grammar is a regular grammar that associates with each derivation a weight, i.e. a vector in the additive monoid $\mathbb{N}^{d}$. Then a derivation is considered successful if the corresponding weight belongs to the monoid generated by $\bf 1^{d}$ $=(1,1,\ldots,1)$. Here the main crux is the construction of a formula (see Proposition 7) based on the computation of the Zariski closure of the intersection of two finitely generated monoids of matrices. The intrinsic difficulty to control the accumulation points of the set – i.e., to exclude those one not coming from the effective functioning of the machines – makes such a computation a non trivial task.
8 Technical details and Examples
We rerun the proofs of Propositions 6 and 11 on the Example 1 of Sec. 4.2.
Example 1 (continued) Let $\cal L$ be the language over the alphabet $\{a,b\}$ given by
=${\cal L}=\{a^{n}b^{n}:n\in\mathbb{N}\}^{*}\ =\ \{a^{n_{1}}b^{n_{1}}\cdots a^{n_{p}}b^{n_{p}}\ :\ n_{1},\ldots,n_{p},\ p\in\mathbb{N}\}.$
Since $\cal L=$ $L^{*},$ with $L=\{a^{n}b^{n}:n\in\mathbb{N}\}$, then
$\mathbf{C}l{(\varphi({\cal L}))}=\mathbf{C}l{(\varphi({L}^{*}))}=\mathbf{C}l{(\varphi({L})^{*})},$
where $\varphi$ is the morphism (4) of the quantum automaton. Recall that $\mathbf{C}l{({\cal A}))}$ and $\overline{{\cal A}}$ denote, respectively, the Euclidean and the Zariski closure of a set $\cal A$. By Theorems 2 and 3, one has $\mathbf{C}l{(\varphi({L})^{*})}=\overline{\varphi({L})^{*}}$. This implies $\mathbf{C}l{(\varphi({\cal L}))}=\overline{\Psi({G})^{*}},$ where
  • $G=M^{*}$ is the monoid generated by the matrix $M\in O_{n}\oplus O_{n}$
    $M=\varphi(a)\oplus\varphi(b)=\left(\begin{array}[]{cc}\varphi(a)&0\\ 0&\varphi(b)\end{array}\right)$
  • $\Psi$ is the regular map defined as $\Psi(A\oplus B):=AB$.
We now prove that, if $\overline{G}$ is irreducible and effective algebraic, then $\overline{\Psi({G})^{*}}$ is in $(\cal H)$ and thus $\mathbf{C}l{(\varphi({\cal L}))}$ is effective algebraic as well.
For this purpose, observe first that $\Psi:O_{n}\oplus O_{n}\rightarrow O_{n}$ is a regular and thus a continuous map. This implies that
$\Psi(\overline{G})\subseteq\overline{\Psi(G)},$
which, in turn, implies
$\overline{\Psi(\overline{G})}\ =\ \overline{\Psi(G)}.$ (44)
Let $H_{1}=\overline{\Psi(\overline{G})}$. By Lemma 1, $H_{1}$ is an algebraic irreducible set. Moreover, being $G$ a group, one has $I\in\Psi(G)$ so $I\in H_{1}$. Since $G$ is finitely generated as a monoid, then, by using Derksen’s algorithm, one can effectively calculate $\overline{G}$. Afterwards, following [12], Sec. 3.1, by using Gröbner bases techniques, since $\overline{G}$ is effective algebraic, then one can effectively compute $H_{1}$.
Let $H_{2}=\overline{H_{1}\cdot H_{1}}$. Observe first that $I\in H_{1}$ implies $I\in H_{2}$. Similarly to the previous case, one observes that, for arbitrary subsets ${\cal A,B}$ of $O_{n}$
[…]
By using the very same argument for $H_{1}$ one verifies that $H_{2}$ is an irreducible effective algebraic set. Moreover, by (44) and (45), one has $H_{2}=\overline{\Psi(G)^{2}}$.
Finally, let us define recursively the family $\{H_{i}\}_{i\geq 1}$ of sets of $O_{n}$, where $H_{0}=\{I\}$ and, for every $i\geq 1$, $H_{i}=\overline{H_{i-1}\cdot H_{1}}.$ Getting an ascending chain
$H_{1}\;\subseteq\;H_{2}\;\subseteq\;\cdots\;H_{i}\;\subseteq\;\cdots$
of irreducible, effective algebraic sets, then there exists some $m\in\mathbb{N}$ such that
[…]
Finally let ${\cal X}=\bigcup_{i\geq 1}\ H_{i}$. By (46), $\overline{\cal X}=H_{1}\cup\cdots\cup H_{m}$. On the other hand, since, for every $i\geq 1$, $H_{i}=\overline{\Psi(G)^{i}}$, then it is checked that
$\overline{\cal X}=H_{1}\cup\cdots\cup H_{m}=\overline{\Psi(G)^{*}}.$
Hence $\overline{\Psi(G)^{*}}$ is effective algebraic.

### v3

@@ 6.2 The case of monoidal languages
By Lemma 7, there exists an effective computable finite subset $\cal R$ of $\ \Sigma_{1}^{*}\times\Sigma_{1}^{*}$ such that $M_{S}=\zeta({\cal R}^{*})$. First let us prove the following claim:
Claim. Let $\cal R$ =$=\{r\}$, with $r\in\Sigma_{1}^{*}\times\Sigma_{1}^{*}$. Then ${\mathbf{C}l}(\zeta({\cal R}))$ contains $I$ and ${\mathbf{C}l}(\zeta({\cal R}))=\Psi(H)$ for some $H\in({\cal H})$ (cf (33)) and regular map $\Psi$.
Let us write $r=(\alpha,\beta)$ with $\alpha=A_{1}A_{2}\cdots A_{s},$ $\beta=B_{1}\cdots B_{t}$, $A_{i},B_{i}\in\Sigma_{1},\ s,t\geq 0$. We assume first that $\alpha$ and $\beta$ have the same length, i.e., $s=t$. By (34), the fact that $\zeta$ is a monoid morphism and Corollary 1, one gets
@@
Since $\zeta$ is a monoid morphism, then one has
${\mathbf{C}l}(M_{S})={\mathbf{C}l}(\zeta({\cal R}^{*}))={\mathbf{C}l}(\zeta({\cal R})^{*}).$
@@
one has ${\mathbf{C}l}(M_{S})={\mathbf{C}l}({\mathbf{C}l}(\zeta({\cal R}))^{*}).$ By Theorems 2 and 3, one has
${\mathbf{C}l}({\mathbf{C}l}(\zeta({\cal R}))^{*})=\overline{{\mathbf{C}l}(\zeta({\cal R}))^{*}},$ (40)
where $\overline{{\mathbf{C}l}(\zeta({\cal R}))^{*}}$ is the Zariski closure of $({\mathbf{C}l}(\zeta({\cal R})))^{*}$. Set =${\cal R}=\{r_{1},\ldots,r_{s}\},$ with $r_{i}\in\Sigma_{1}^{*}\times\Sigma_{1}^{*},s\geq 1$. Since, for every $i=1,\ldots,s$, ${\mathbf{C}l}(\zeta({r_{i}}))$ contains the identity, by Lemma 2, one gets ${\mathbf{C}l}(\zeta({\cal R}))^{*}=({\mathbf{C}l}(\zeta({r_{1}}))\cdots{\mathbf{C}l}(\zeta({r_{s}})))^{*}$, which, thus, implies
${\mathbf{C}l}(M_{S})=\overline{({\mathbf{C}l}(\zeta({r_{1}}))\cdots{\mathbf{C}l}(\zeta({r_{s}})))^{*}}.$
By the claim, for every $i=1,\ldots,s$, ${\mathbf{C}l}(\zeta({r_{i}}))=\Psi(H_{i})$, where $H_{i}\in({\cal H})$, and $\Psi$ is a regular map not depending upon $H_{i}$. Hence one gets
${\mathbf{C}l}(M_{S})\ =\ \overline{(\Psi(H_{1})\ \cdots\ \Psi(H_{k}))^{*}}.$ (41)
Observe that in (41) we can write $\Psi(H_{1})\ \cdots\ \Psi(H_{k})$ as $\widehat{\Psi}(H_{1}\oplus\cdots\oplus H_{k})$, where $\widehat{\Psi}$ is the regular map defined as $\widehat{\Psi}(X_{1}\oplus\cdots\oplus X_{k})=\Psi(X_{1})\cdots\Psi(X_{k})$. Hence we get
${\mathbf{C}l}(M_{S})=\overline{(\widehat{\Psi}(H))^{*}},$
where, again, $H=(H_{1}\oplus\cdots\oplus H_{k})\in({\cal H})$ as one can easily check. Finally ${\mathbf{C}l}(M_{S})\in({\cal H})$ follows by applying Proposition 6. ∎∎
@@
Proof of Proposition 11: As remarked before, it is enough to prove the claim for $k\leq 2$ since, by (19), the general case is treated with the same argument. Set $G=\langle V,\Sigma,P,S\rangle$ and observe that, for every $A\in V$, the fact that ${\mathbf{C}l}(M_{A})$ is effective algebraic comes from Lemmata 4 and 8. The claim now follows from Proposition 9. ∎
@@ 7 Concluding remarks and open problems
[$\alpha$]  One could extend Proposition 11 to a broader class of quantum automata. In this context, the construction of the decidability procedure would require the computation of the Zariski closure of a group $G$ of matrices that is algebraically presented, i.e. $G$ admits an infinite set of generators that is algebraic itself. It may be of interest to investigate conditions ensuring the extension of the algorithms of [15, 23, 35] for such groups.
@@
[$\gamma$]  The $(L,Q)$ Intersection problem is decidable for bounded semi-linear languages. It may be of interest to get extensions of this result for the broader class of languages accepted by the reversal bounded non deterministic counter machines [25]. This extension could exploit a characterization of such languages provided by Ibarra in [26] in terms of some special combinatorial systems called RLGMC-grammars (Right-Linear Grammar with Multi-Counters). A RLGMC-grammar is a regular grammar that associates with each derivation a weight, i.e. a vector in the additive monoid $\mathbb{N}^{d}$. Then a derivation is considered successful if the corresponding weight belongs to the monoid generated by $\bf 1^{d}$ $=(1,1,\ldots,1)$. In this case, the construction of the formula that makes the problem decidable, is reduced to the computation of the Zariski closure of the intersection of two finitely generated monoids of matrices, say $M_{1}$ and $M_{2}$. One can compute $\overline{M_{1}}$, $\overline{M_{2}}$ and, consequently, $\overline{M_{1}}\cap\overline{M_{2}}$. However, in general, the inclusion $\overline{M_{1}\cap M_{2}}\subseteq\overline{M_{1}}\cap\overline{M_{2}}$ is strict and the intrinsic difficulty to control the accumulation points of the set – i.e., to exclude those not coming from the effective functioning of the machines – makes such a computation a non trivial task.
Acknowledgments The authors would like to thank Mima Stanojkovski for useful discussion on her paper [16].
Appendix
We recall the construction of the context-free grammar that generates the image $\theta(L)$ of a context-free language $L$ over the alphabet $\Sigma_{1}$ under a context-free substitution $\theta:\Sigma_{1}^{*}\rightarrow\Pi(\Sigma_{2}^{*})$, i.e. a $\cal F$-substitution, with $\cal F$=CFL.
We follow verbatim the argument of [6], Sec. II.2. Let ${\cal G}_{1}=\langle V_{1},\Sigma_{1},P_{1},S\rangle$ be the context-free grammar that generates $L$ and let
${\cal G}_{2}=\{G_{x}\ :\ x\in\Sigma_{1}\}$
be the family of context-free grammars where, for every $x\in\Sigma_{1}$,
$G_{x}=\langle V_{x},\Sigma_{2},P_{x},S_{x}\rangle$
is the grammar that generates $\theta(x)$. One may assume that, for all $x\in\Sigma_{1}$, $V_{1}\cap V_{x}=\emptyset$ and, for every $x,y\in\Sigma_{1},\ x\neq y$, $V_{x}\cap V_{y}=\emptyset$. Let be the set
${\cal V}_{2}=\{S_{x}:x\in\Sigma_{1}\}$
and let the copy morphism
$\mbox{c}:(V_{1}\cup\Sigma_{1})^{*}\rightarrow(V_{1}\cup{\cal V}_{2})^{*},$ (42)
$G_{\mbox{c}}=\langle V_{1},\{S_{x}:x\in\Sigma_{1}\},P_{\mbox{c}},S\rangle,$
where $P_{\mbox{c}}=\{A\rightarrow\mbox{c}(\alpha):(A\rightarrow\alpha)\in P_{1}\}$. Finally, define
$H=\langle W,\Sigma_{2},Q,S\rangle$ (43)
as the context-free grammar where
Such grammar $H$ is called the composition of ${\cal G}_{1}$ and ${\cal G}_{2}$ and denoted by
$G={\cal G}_{1}\circ{\cal G}_{2}.$ (44)
The following lemma holds.
Lemma 9
If $G$ is the grammar (44), then $\theta(L)=L({G})$.
A composition of $k$ families of context-free grammars $({\cal G}_{1},\ldots,{\cal G}_{k}),$ with $k\geq 1$ will be the context-free grammar
$G={\cal G}_{1}\circ{\cal G}_{2}\circ\cdots\circ{\cal G}_{k},$
obtained by iterating $(k-1)$ times the operation (44) of composition of grammars on the families ${\cal G}_{1},\ldots,{\cal G}_{k}$.